OpenStudy (anonymous):

The area of a trapezoid is 144 ft 2. If the height is 8 feet and one of the bases is 12 feet, what is the length of the other base? 7.2 ft 14.4 ft 24 ft 16 ft

OpenStudy (anonymous):

the formula for the area of trapezoid is going to be (B1+B2)h / 2 144 = (12+x)8 /2 288 = (12+x)8 36 = 12+x x =24

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

No prob